OpenStudy (anonymous):

write the equation for the following graph (-5,-8),(0,-5),(0,8)

OpenStudy (anonymous):

does anyone know what this means?

ganeshie8 (ganeshie8):

think of it like this : 1) between (-5, 0) it has a slope of 3/5 2) at 0 its value is 8 can you write equation now ?

OpenStudy (anonymous):

That is not a line, I think. It may be \[ x=-\frac{5 y^2}{48}+\frac{5 y}{16}+\frac{25}{6} \] Stay tuned to explain it to you.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

You write \[ f[y] = a y^2 + b y + c\\ f[-8] = -5 \\ f[-5] = 0 \\ f[8] = 0 \] and you solve the three equations for a, b and c

OpenStudy (anonymous):

you get \[ \begin{array}{ccc} a=-\frac{5}{48}, & b=\frac{5}{16}, & c=\frac{25}{6} \\ \end{array} \]

OpenStudy (anonymous):

There is no one line going through the three points. There is a parabola. This is the one I found above.

ganeshie8 (ganeshie8):

this is just a line & a point 1) y+8 = (3/5)(x+5) {-5<= x < 0} 2) y = 8 {x = 0}

OpenStudy (anonymous):

ok I got what you are saying im trying to look now and make sure i wrote this question down correct.

OpenStudy (anonymous):

If the question is ti find one equation to fit the three points, then it is a parabola The one I found.

OpenStudy (anonymous):

See the graph going through your red points.

OpenStudy (anonymous):

im sorry my school page is just taking so long to load

OpenStudy (anonymous):

Did you see the graph?